Help me solve #21 please what numbers do you use
ing balance sheet and income eeded for the next three questions: The following statement for Central Manufacturing Corporation Balance Sheet Assets: Cash Accounts Receivable Inventories Land Other Fixed Assets Income Statement $1,500 8,000 7,500 15,000 14,000 Sales (all credit) $57,500 Cost of Goods Sold 33,5001 24,000 Operating Expenses (20,000) Depreciation iabilities &Owners Equity Accounts Payable Long Term Debt Common Stock Retained Earnings 3,800 5,400 30,000 6,800 3,000 Interest Expense(450 2,550 8501 Taxes What is Centrals net profit margin? 19. a) 4.43% c) 3.7% d) 7,08% e))2.96% b) 5.22% What is Centrals return on equity? a) 4.03% 20. NET iniom c) 6.93% ) 4.62% e) 3.7% b) 5.67% ban n hum Cost of Inu 21. Calculate Centrals operating income return on investment. e) 3.7% a) 6.52% b) 52.17% c) 5.54% d) 8.7% 22 Which of the following is true regarding the over-the-counter market? a) The OTC market is not regulated by the SEC. b) The OTC market includes most security exchanges. Q) The OTC brokers and dealers are linked by NASDAQ. d) The OTC market is a primary market. e) All of the above are true. of securities is marketed to a definite and select group of inv louees or current stockholders, the issue is called a(n): w issue When a ne
Answer #1

21. Operating Income return on investment = Operating Income/investment

Operating income = Net profit + taxes + interest expense = 1700+850+450 = 3000

Therefore, Operating income ROI = 3000/46000 (total assets) = 0.0652 or 6.52%